A proton is in moving in a circular orbit radius 14cm I a uniform 0.35T magnetic field perpendicular to the velocity and proton find speed of proton


1
Expert's answer
2021-01-19T07:08:17-0500

Given:

"R=0.14\:\rm m"

"B=0.35\:\rm T"

"m=1.67\times10^{-27}\:\rm kg"

"q=1.60\times10^{-19}\:\rm C"

The Newton's second law says


"\frac{mv^2}{R}=qvB"

Hence


"v=\frac{qBR}{m}\\
v=\frac{1.60\times 10^{-19}\times 0.35\times 0.14}{1.67\times10^{-27}}=4.7\times 10^6\:\rm m/s"